OpenStudy (ksaimouli):

bailey decorates dance hall with balloons.A box of balloons contains 20 red,18blue,and 12 white balloons.bailey randomely choose a balloon without looking at it.what is probability that he choose ablue or white

OpenStudy (mathmate):

A blue or white means not red. Total number of balloons = 20+18+12=50 P(red)=20/50=0.4 P(~red)=1-0.4=0.6 Alternatively, P(blue or white)=P(blue)+P(white)=18/50+12/50=0.6

OpenStudy (ksaimouli):

how did u get .1-.4